OpenStudy (anonymous):

What number must be subtracted from -12 to obtain 24 ? What number must be subtracted from -20 to obtain 15 ?

OpenStudy (kkutie7):

-12-x=24 -x=36 x=-36 -20-x=15 -x=35 x=-35

OpenStudy (anonymous):

so its -36 and -35 @Kkutie7

OpenStudy (kkutie7):

check the work. -12-(-36)= -12+36=24 and -20-(-35)=15 so yes

OpenStudy (kkutie7):

OK good =)